pV e 4 o, o D / İ + 0.CAY-TA dau at Nou wad to nov gOo n coloe much do uu to pou tby do Dhilds collhoe
Answer #1

Answer for 1st question

What will 110000 grow to be in 9 years if its invested today at 11%

= 110000(1+0.11)^9

= 110000(2.558037)

= 281384.0617

Answer for 2nd question

You would like to have 200000 in college funds in 15 years. How much would you need to invest if you expect to earn 12%...

FV = PV(1+r)^n

200000 = PV(1+0.12)^15

200000 = PV(5.47356576)

PV = 200000/5.47356576

Therefore amounted needed to be invested = 36539.252253